OpenStudy (anonymous):

The line of symmetry for the quadratic equation y = ax^2 + 8x - 3 is x = 4. What is the value of "a"? -2?

OpenStudy (anonymous):

a determines the "shape", the steepness of the curve. OK?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

The line of symmetry is (I believe) the line that splits the parabola in two equal parts, here the vertical line that goes through the vertex. OK?

OpenStudy (anonymous):

So the question is, do you think that knowing the steepness of the curve can tell you anything about its vertex, or alternatively, knowing the vertex can tell you anything about the steepness?

OpenStudy (anonymous):

OH.. screwed up, of course

OpenStudy (anonymous):

vertex is at x = -b/2a = 4 b = 8

OpenStudy (anonymous):

?? i dun get it can you explain it more easier plz

OpenStudy (anonymous):

There is an equation you need to memorize. At the vertex (and along the line of symmetry), x = -b/2a.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

We're given x and know b from the equation, so we can compute a.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

We have x = 4 = -b/2a = -8/(2a) = -4/a

OpenStudy (anonymous):

y=ax^2+8x-3 y=a4^2+32-3 y=a4^2+29 y=(1)4^2+29

OpenStudy (anonymous):

Huh? a = -1 Check by y = -x^2 + 8x -3 vertex is at -b/2a = -8/2(-1) = 4 check

OpenStudy (anonymous):

oh y=(-1)4^2+29

OpenStudy (anonymous):

y=13

OpenStudy (anonymous):

Good. But the question only asks for a.